### v1

@@ 1 Introduction
Perhaps the most classic example of this approach is the scale that was developed by Kardashev (1964) for describing “technologically developed civilizations” as they expand through space. The underlying assumption of the Kardashev (1964) is that the growth in power consumed by human civilization—and perhaps a technological civilization in general—will continue to the point where expansion into space becomes necessary:
@@
This model of continuous growth for thousands of years into the future has been used to motivate observational and theoretical approaches to technosignature science by searching for civilizations that might follow such trajectories. For example, Sagan (1973) suggested that “the best policy might therefore be to search with existing technology for Type II or Type III civilizations among the nearer galaxies, rather than Type I or younger civilizations among the nearer stars.” The analysis by Wright et al. (2022) similarly asserted that “technosignatures, by contrast [to other biosignatures], have essentially no upper limit to their detectability,” with a reference to the Kardashev (1964) scale. Others have critiqued the idea of continuous growth and suggested that any long-lived civilizations might instead expand at a slower rate, or not expand into space at all (e.g., Von Hoerner, 1975; Newman and Sagan, 1981; Haqq-Misra and Baum, 2009; Mullan and Haqq-Misra, 2019). Such arguments, and their rebuttals, have all drawn upon informal methods that contain implicit assumptions, which inevitably biases any results toward what is familiar and known.
The interdisciplinary field of futures studies has developed a variety of systematic methodological approaches toward developing self-consistent future trajectories. The field uses the plural “futures” to indicate that it does not attempt to predict a singular future but instead is an approach to project multiple different futures that illustrate a range of possible scenarios. Such methods attempt to avoid both technological determinism and social determinism when discussing future technological developments; instead, human technology and human society are seen as intertwined, interacting, co-evolving, and capable of mutually shaping each other. Rather than making reductionist assumptions about continual rates of growth (e.g., Kardashev, 1964), futures studies methods draw upon insights from social sciences and other disciplines to understand the intertwined influences of human cultures with their political, economic, and social systems.
@@
Technosignature science can benefit from a more systematized approach toward envisioning possible futures, with trajectories of Earth’s futures serving as self-consistent scenarios for the assessment of technosignature detectablility and search strategy. The use of futures studies methods has been suggested by Voros (2018) as a relevant tool for mapping the morphology of extraterrestrial contact scenarios, while Profitiliotis and Theologou (2023) have suggested that training in futures studies methods and ways of thinking (known as “futures literacy”) could help to prepare for the unknown dimensions of discovering extraterrestrial life, including intelligent life. In this study, we use scenarios of Earth’s future as templates for possible extraterrestrial technology. The greatest difficulty in applying futures studies methods to the projection of Earth’s future technosignatures is the timescale: existing methods in futures studies make projections on short timescales of years to decades, but no generalized methods exist for envisioning futures on much longer century to millennium timescales. The development of detailed projections across long time scales is difficult, but the problem becomes more tractable when the goal is to predict changes in Earth’s technosphere that could be detectable through astronomical observation or exploration.
@@ 2 Scenario Modeling
The general morphological analysis has three steps, which we briefly describe before demonstrating the process. In step one, we first state the problem in as exact a formulation as possible, and then break the problem down into a number of parameters that are “meaningful, equally important, abstract, straightforward, independent of each other, and have many internal connections” (Johansen, 2018). In step two, we then use these parameters to construct a multidimensional matrix that describes the full scenario space, where each unique combination of parameter values defines a scenario. In step three, we perform a cross-consistency assessment to eliminate any inconsistent scenarios (i.e., those that are logically impossible or paradoxical), which will result in a final set of plausible and self-consistent scenarios for further analysis.
@@
Here we state the problem as: “What are the technological phenomena of the future anthroposphere and how can they be described?” The phrasing of this question draws upon Earth system science (Steffen et al., 2020), with the anthroposphere as part of the larger Earth system. The technosphere is included as a subset of the anthroposphere, along with other subsystems such as economy, institutions, and cultures. We can consider the technosphere as the material manifestation of infrastructure that is shaped by the immaterial subsystems of the anthroposphere. This statement of the problem enables our analysis to focus on understanding the technological phenomena that emerge from the plausible set of future anthropospheres. It’s worth noting that our analysis focuses only on anthropospheres that are viable and functional as continuities when “observed,” which makes them relevant and interesting under a principle of mediocrity. In other words, we do not investigate anthropospheres that have completely collapsed to the point of human extinction or are on the verge of particular turning points that are effectively special occasions of discontinuity in the duration of 1000-year civilizational timescales.
@@
We break down the problem using the widely applied PEST (political, economic, social, and technological) framework (Aguilar, 1967), which provides four factors that can provide meaningful and important descriptions of future scenarios. Our interest in this study is specifically in understanding the possibility space for Earth’s future technosphere, so we further decompose our parameter space into two sets: one focusing on political, economic, and social factors (§2.1) and another focusing specifically on technology factors (§2.2). For each of these two sets, we develop a multidimensional matrix based on a fixed number of parameters, each of which can be assigned to one of several possible values in an exhaustive way. We then assess each matrix for cross-consistency and develop a final set of scenarios based on the remaining sets of factors from the two matrices.
@@ 2.1 Global Factors
We start the general morphological analysis by focusing on the first three factors of the PEST framework: the political, economic, and social factors. We use these three factors to define our first morphological matrix, which describes future human society in terms of its global institutions. The purpose of this study is to characterize detectable properties of planetary technospheres, so our assignment of political, economic, and social factors will be representative of global trends taken in aggregate. This approach does not account for any local or regional variations at this stage, but such heterogeneities will be included later during the worldbuilding phase (§3). We construct the global factors morphological matrix in Table 1 with these three parameters to describe the global economic, political, and social systems for our scenario space. These factors are further described in this subsection.
@@ X Economy Y Politics C Society
The first factor is the global economic system (X). This factor addresses the issue of producing and distributing scarce resources among groups/communities of humans. Specific solutions for economic systems are numerous, but we attempt to avoid restricting our long-term future projections to any specific known solutions. We instead consider the possibilities that the future economy will operate either under scarcity (X1, most resources are limited) or achieve non-scarcity (X2, most resources are unlimited). A scarcity economy is consistent with present-day Earth, in which nearly all commodities are limited and prices can vary based on availability. A non-scarcity economy, by contrast, would have most or all resources available as unlimited, especially for basic human survival needs, in the way that breathable air is generally not considered as a commodity on Earth today.
@@
The second factor is the global political system (Y). This factor addresses the issue of rule or sovereignty over the collective lives of groups, states, or other associations of humans. The choice of values for this factor should also avoid the tendency to rely on any existing known political theories and instead should take general form that can be extended to long-term futures. We consider four options for the global political system to capture a complete set of possibilities. Rule by one (Y1) involves a single actor with a complete hold of global political power, rule by few (Y2) involves a small number of actors that hold the majority of global political power, rule by all (Y3) involves global political power being widely disseminated, and rule by none (Y4) involves the absence of or a breakdown in the organization of global politics. The Y2 factor is the most consistent with present-day Earth.
@@
The third factor is the global social system (Z), which addresses the issue of how groups, states, or other associations of humans are organized. As with the other factors, the values for this factor should take a general form to be applicable to long-term scenarios. We select two possible options for the global social system, which can be either organization in hierarchical structures (Z1) in which top-down social dynamics dominate or distributed structures (Z2) in which horizontal social interactions are primary. The Z1 factor is the most consistent with present-day Earth.
@@
The general approach to a cross-consistency assessment is to inspect each pair of values in the multidimensional matrix to determine whether a contraction or inconsistency exists. There are twenty unique possible combinations of value pairs in Table 1. For this study, we use Anthropic’s “Claude,” [^1] https://www.anthropic.com/claude a large language model (LLM) that has been trained using the “Constitutional AI” process (Bai et al., 2022), which sets it apart from other LLMs. This process trained Claude to be helpful, honest, and harmless, not via human feedback labels for harmful outputs but via self-improvement steered by a small set of principles that comprise a “constitution” that governs its behavior. Furthermore, Artificial Intelligence-assisted scenario development based on LLMs is already being explored within the futures studies community and has been described as “promising” for the generation of base material to be considered further by human experts in a hybrid strategic foresight process (Spaniol and Rowland, 2023). Combining the systematized set of modeled scenarios constructed via general morphological analysis with the generative strengths of Claude, which was found by the authors to be competent in scenario development methods, presents an excellent opportunity for producing an initial substrate for evaluation and modification by human experts. We harness this opportunity by invoking the LLM to conduct a first cross-consistency assessment of the global factors value pairs with explicit concise arguments for and against cross-consistency, which we then review manually and correct as needed. (We note that the use of the LLM saved significant time but still resulted in some erroneous assessments. The assessment was facilitated by the LLM’s arguments, but was eventually conducted by humans, which is the standard practice in general morphological analysis.) The results of this cross-consistency assessment are shown in Figure 1, where the cells marked with an ‘x’ are those that have been deemed inconsistent. This gives a cross-consistency matrix in which eight of the twenty possible value pairs have been eliminated.
@@
Figure 1: Global factors cross-consistency assessment. The 8 cells marked ‘x’ indicate inconsistent value pairs. The remaining 12 open cells indicate consistent value pairs.
@@
With the cross-consistency assessment complete, the remaining value pairs in the global factors cross-consistency matrix (Fig. 1) result in three viable scenarios, which are summarized in Table 2. To better understand the features and differences provided by each of these combinations, we use the Claude LLM to generate illustrative descriptions of these three global factors scenarios for a thousand years into the future. The generated text for each scenario provides an initial basis as an entry point for our deeper (human) analysis of the themes underlying each scenario. We assign an “archetype” to each scenario based on the four generic images of the future system created by Dator (see., e.g., Dator, 2009; Bezold, 2009): in Dator’s general model, any given scenario can be characterized as undergoing a state of growth (upward trajectory), a state of collapse (downward trajectory), a state of discipline (steady or horizontal trajectory), or a state of transformation (non-linear or disjointed trajectory). We also assign a “myth/metaphor” for each scenario to reflect its guiding civilizational undercurrent, which draws upon a method known as “causal layered analysis” (Inayatullah, 1998) applied to understand the deepest underlying motivations of the actors and systems in a given scenario. In our case, we critically evaluate the illustrative descriptions generated by the LLM to identify the most salient way of describing the themes that emerge from each scenario.

### v2

@@ 1 Introduction
Perhaps the most classic example of this approach is the scale that was developed by Kardashev (1964) for describing “technologically developed civilizations” as they expand through space. We focus specifically on the Kardashev scale because this idea has become foundational in the technosignature research community (see, e.g., Sagan, 1973; Carrigan Jr, 2012; Cirkovic, 2015; Besteiro, 2019; Wright et al., 2022; Gray, 2020) and is also frequently invoked in other research that seeks to project possible trajectories of human civilization (see, e.g., Baker, 2020; Namboodiripad and Nimal, 2021; Jiang et al., 2022; Zhang et al., 2023). The underlying assumption of Kardashev (1964) is that the growth in power consumed by human civilization—and perhaps a technological civilization in general—will continue to the point where expansion into space becomes necessary:
@@
This model of continuous growth for thousands of years into the future has been used to motivate observational and theoretical approaches to technosignature science by searching for civilizations that might follow such trajectories. For example, Sagan (1973) suggested that “the best policy might therefore be to search with existing technology for Type II or Type III civilizations among the nearer galaxies, rather than Type I or younger civilizations among the nearer stars.” The analysis by Wright et al. (2022) similarly asserted that “technosignatures, by contrast [to other biosignatures], have essentially no upper limit to their detectability,” with a reference to the Kardashev (1964) scale. Others have critiqued the idea of continuous growth and suggested that any long-lived civilizations might instead expand at a slower rate, or not expand into space at all (e.g., Von Hoerner, 1975; Newman and Sagan, 1981; Haqq-Misra and Baum, 2009; Mullan and Haqq-Misra, 2019; Likavčan, 2024). Such arguments, and their rebuttals, have all drawn upon informal methods that contain implicit assumptions, which inevitably biases any results toward what is familiar and known.
The interdisciplinary field of futures studies has developed a variety of systematic methodological approaches toward developing self-consistent future trajectories. The field uses the plural “futures” to indicate that it does not attempt to predict a singular future but instead is an approach to project multiple different futures that illustrate a range of possible scenarios. Such methods attempt to avoid both technological determinism and social determinism when discussing future technological developments; instead, human technology and human society are seen as intertwined, interacting, co-evolving, and capable of mutually shaping each other. Rather than making reductionist assumptions about continual rates of growth, futures studies methods draw upon insights from social sciences and other disciplines to understand the intertwined influences of human cultures with their political, economic, and social systems.
@@
Technosignature science can benefit from a more systematized approach toward envisioning possible futures, with trajectories of Earth’s futures serving as self-consistent scenarios for the assessment of technosignature detectablility and search strategy. The use of futures studies methods has been suggested by Voros (2018) as a relevant tool for mapping the morphology of extraterrestrial contact scenarios, while Profitiliotis and Theologou (2023) have suggested that training in futures studies methods and ways of thinking (known as “futures literacy”) could help to prepare for the unknown dimensions of discovering extraterrestrial life, including intelligent life. The use of scenarios has also been suggested by Ramirez et al. (2015) as a rigorous methodology for generating “interesting research” that can enable scholars to meaningfully engage with complex and uncertain problems. In this study, we use scenarios of Earth’s future as templates for possible extraterrestrial technology. The greatest difficulty in applying futures studies methods to the projection of Earth’s future technosignatures is the timescale: existing methods in futures studies make projections on short timescales of years to decades, but no generalized methods exist for envisioning futures on much longer century to millennium timescales. The development of detailed projections across long time scales is difficult, but the problem becomes more tractable when the goal is to predict changes in Earth’s technosphere that could be detectable through astronomical observation or exploration.
@@ 2 Scenario Modeling
The general morphological analysis has three steps, which we briefly describe before demonstrating the process. In step one, we first state the problem in as exact a formulation as possible, and then break the problem down into a number of parameters that are “meaningful, equally important, abstract, straightforward, independent of each other, and have many internal connections” (Johansen, 2018). In step two, we then use these parameters to construct a multidimensional matrix that describes the full scenario space, where each unique combination of parameter values defines a scenario. In step three, we perform a cross-consistency assessment to eliminate any inconsistent scenarios (i.e., those that are logically impossible or paradoxical), which will result in a final set of plausible and self-consistent scenarios for further analysis. We note that our use of the term “self-consistent” refers to our use of a cross-consistency assessment (described further below) to eliminate scenarios that would contain self-contradictory, impossible, or unlikely qualities, so that our final set of scenarios are only those that represent plausible and realizable futures.
@@
Here we state the problem as: “What are the technological phenomena of the future anthroposphere and how can they be described?” The phrasing of this question draws upon Earth system science (Steffen et al., 2020), with the anthroposphere as part of the larger Earth system. The technosphere is included as a subset of the anthroposphere, along with other subsystems such as economy, institutions, and cultures. We can consider the technosphere as the material manifestation of infrastructure that is shaped by the immaterial subsystems of the anthroposphere. This statement of the problem enables our analysis to focus on understanding the technological phenomena that emerge from the plausible set of future anthropospheres. It’s worth noting that our analysis focuses only on anthropospheres that are viable and functional as continuities when “observed,” which makes them relevant and interesting under a principle of mediocrity. In other words, we do not investigate anthropospheres that have completely collapsed to the point of human extinction or are on the verge of particular turning points that are effectively special occasions of discontinuity in the duration of 1000-year civilizational timescales. We acknowledge the possibility that “graveyard technosignatures” (see, e.g., Haqq-Misra and Baum, 2009; Villarroel et al., 2022) that serve as lingering evidence of a now-extinct technological civilization are a possible subject of interest and even could be the most prominent technosignature in the galaxy or universe. However, including graveyard systems is beyond the scope of our analysis, which is based on the coupling between human and technological systems. We acknowledge that further extensions of our methods could be applied to graveyard scenarios, but we leave such considerations for a subsequent study.
@@
We break down the problem using the widely applied PEST (political, economic, social, and technological) framework (Aguilar, 1967), which provides four factors that can provide meaningful and important descriptions of future scenarios. We note that many variations of the PEST framework exist, which can include explicit consideration of environmental factors as well as legal factors (such as the PESTEL framework). Our approach does not neglect environmental (or legal) implications, but we do not prescribe these factors during our scenario modeling. Instead, environmental (and relevant legal) factors are all emergent properties of our scenarios, which result from the political, economic, social, and technological factors that define each scenario. Our interest in this study is specifically in understanding the possibility space for Earth’s future technosphere, so we further decompose our parameter space into two sets: one focusing on political, economic, and social factors (§2.1) and another focusing specifically on technology factors (§2.2). For each of these two sets, we develop a multidimensional matrix based on a fixed number of parameters, each of which can be assigned to one of several possible values in an exhaustive way. We then assess each matrix for cross-consistency and develop a final set of scenarios based on the remaining sets of factors from the two matrices. We note at the outset that our approach is intended to capture the interdependencies that exist between political, economic, social, and technological factors, rather than assuming that any of these is necessarily determined exclusively by others. Our use of a second matrix for technology factors alone does not at all imply that technology is determined by political, economic, or social factors (or vice versa) and instead should be interpreted as a bidirectional relationship of co-shaping. As will be seen in the approach that follows, our separate selection of global factors and technological factors is to enable an assessment of bidirectional consistency between these two sets of factors, rather than implying a deterministic dependency of one on the other.
@@ 2.1 Global Factors
We start the general morphological analysis by focusing on the first three factors of the PEST framework: the political, economic, and social factors. We use these three factors to define our first morphological matrix, which describes future human society in terms of its global institutions. The purpose of this study is to characterize detectable properties of planetary technospheres, so our assignment of political, economic, and social factors will be representative of global trends taken in aggregate. This approach does not account for any local or regional variations at this stage, but such heterogeneities will be included later during the worldbuilding phase (§3). We construct the global factors morphological matrix in Table 1 with these three parameters to describe the global economic, political, and social systems for our scenario space. We selected the particular values for each of these factors deductively based on our need to be both generalizable and comprehensive in the construction of our matrix of the immaterial part of the anthroposphere. A generalized set of values for these parameters is necessary for making projections on a 1000-year time horizon because the set of present-day political, economic, and social ideologies and philosophies will not necessarily reflect the set of such ideologies that are present in any given future; rather than employing horizon scanning to identify trends, uncertainties, and weak signals that may have inductively helped map out multiple alternatives for the evolution of current ideologies over time horizons of two to three decades, we have adopted a more generalized set of values that can capture the essence of both existing ideologies and any new possibilities that might emerge later in the course of our 1000-year time horizon. A comprehensive set of values refers to our choice of the most fundamental units of variability that can logically occur in the possibility space for each parameter; such choices are intended to construct our matrix in a way that includes all permutations of how political, social, and economic factors could change in the most fundamental ways. These factors are further described in this subsection.
@@ X Economy Y Politics C Society
The first factor is the global economic system (X). This factor addresses the issue of producing and distributing scarce resources among groups/communities of humans. Specific solutions for economic systems are numerous, but we attempt to avoid restricting our long-term future projections to any specific known solutions. We instead consider the possibilities that the future economy will operate either under scarcity (X1, most resources are limited) or achieve non-scarcity (X2, most resources are unlimited). A scarcity economy is consistent with present-day Earth, in which nearly all commodities are limited and prices can vary based on availability. A non-scarcity economy, by contrast, would have most or all resources available as unlimited, especially for basic human survival needs, in the way that breathable air is generally not considered as a commodity on Earth today. These values for the economy factor are admittedly a simplification of the range of existing economic ideologies, but the use of a binary selection between scarcity and non-scarcity provides a generalizable approach to grossly categorize current and future ideologies based on their most fundamental essence; likewise, these choices allow for a comprehensive description of future economic systems, as the choice between scarcity and non-scarcity possibilities covers the full range of possibilities.
@@
The second factor is the global political system (Y). This factor addresses the issue of rule or sovereignty over the collective lives of groups, states, or other associations of humans. The choice of values for this factor should also avoid the tendency to rely on any existing known political theories and instead should take general form that can be extended to long-term futures. We consider four options for the global political system to capture a complete set of possibilities. Rule by one (Y1) involves a single actor with a complete hold of global political power, rule by few (Y2) involves a small number of actors that hold the majority of global political power, rule by all (Y3) involves global political power being widely disseminated, and rule by none (Y4) involves the absence of or a breakdown in the organization of global politics. The Y2 factor is the most consistent with present-day Earth. As with the economy factor, the values chosen for this political factor are also intended to be generalizable beyond present-day political ideologies; likewise, the selection of four values provides a comprehensive description of the possible configurations of possible political systems at the most fundamental unit of variability.
@@
The third factor is the global social system (Z), which addresses the issue of how groups, states, or other associations of humans are organized. As with the other factors, the values for this factor should take a general form to be applicable to long-term scenarios. We select two possible options for the global social system, which can be either organization in hierarchical structures (Z1) in which top-down social dynamics dominate or distributed structures (Z2) in which horizontal social interactions are primary. The Z1 factor is the most consistent with present-day Earth. The values chosen for the social factor are abstracted in the most generalizable way, to reflect either a vertical or horizontal, i.e., egalitarian, distribution of power as the most prominent feature; this binary selection between hierarchical and distributed systems likewise is comprehensive, with the assumption that any present-day or future social system will be fundamentally characterized by one of these two choices.
@@
The general approach to a cross-consistency assessment is to inspect each pair of values in the multidimensional matrix to determine whether a contraction or inconsistency exists. There are twenty unique possible combinations of value pairs in Table 1. For this study, we use Anthropic’s “Claude,” [^1] https://www.anthropic.com/claude a large language model (LLM) that has been trained using the “Constitutional AI” process (Bai et al., 2022), which sets it apart from other LLMs. This process trained Claude to be helpful, honest, and harmless, not via human feedback labels for harmful outputs but via self-improvement steered by a small set of principles that comprise a “constitution” that governs its behavior. Furthermore, Artificial Intelligence-assisted scenario development based on LLMs is already being explored within the futures studies community and has been described as “promising” for the generation of base material to be considered further by human experts in a hybrid strategic foresight process (Spaniol and Rowland, 2023). Combining the systematized set of modeled scenarios constructed via general morphological analysis with the generative strengths of Claude, which was found by the authors to be competent in scenario development methods, presents an excellent opportunity for producing an initial substrate for evaluation and modification by human experts. It should be noted that we only utilized the LLM as a descriptive tool, prompting it to express the multiple parameter value combinations in natural language and to describe in natural language the ramifications of particular value pairs being present together. We harness this opportunity by invoking the LLM to conduct a first cross-consistency assessment of the global factors value pairs with explicit concise arguments for and against cross-consistency, which we then review manually and correct as needed. (We note that the use of the LLM saved significant time but still resulted in some erroneous assessments. The assessment was facilitated by the LLM’s arguments, but was eventually conducted by humans, which is the standard practice in general morphological analysis.) We use a 4-point Likert scale to evaluate each value pair’s consistency: 0 for no consistency, 1 for low consistency, 2 for moderate consistency, and 3 for high consistency. For our analysis of global factors, we retain value pairs of moderate and high consistency for the next steps of the analysis, while the rest are rejected. The results of this cross-consistency assessment are shown in Figure 1, where the cells with a score of 2 or 3 indicate consistent value pairs and those with a score less than 2 have been deemed inconsistent. This gives a cross-consistency matrix in which eight of the twenty possible value pairs have been eliminated.
@@
Figure 1: Global factors cross-consistency assessment. The 12 green cells with scores of 2 or 3 indicate consistent value pairs. The remaining 8 cells with scores less than 2 indicate inconsistent value pairs.
@@
With the cross-consistency assessment complete, the remaining value pairs in the global factors cross-consistency matrix (Fig. 1) result in three viable scenarios, which are summarized in Table 2. To better understand the features and differences provided by each of these combinations, we prompt the Claude LLM to express in natural language the value combinations of these three global factors scenarios for a thousand years into the future, generating illustrative descriptions. The generated text for each scenario provides an initial basis as an entry point for our deeper (human) analysis of the themes underlying each scenario. Without assistance from the LLM, we deductively assign an “archetype” to each scenario based on the four generic images of the future system created by Dator (see, e.g., Dator, 2009; Bezold, 2009): in Dator’s general model, any given scenario can be characterized as undergoing a state of growth (upward trajectory), a state of collapse (downward trajectory), a state of discipline (steady or horizontal trajectory), or a state of transformation (non-linear or disjointed trajectory). These archetypes help us frame subsequent worldbuilding options after the combination of these intermediate scenarios with the technology factors intermediate scenarios into the complete, final, nuanced anthroposphere scenarios. Next, again without assistance from the LLM, we inductively assign a “myth/metaphor” for each scenario to reflect its guiding civilizational undercurrent, which draws upon a method known as “causal layered analysis” (Inayatullah, 1998) applied to understand the deepest underlying motivations of the actors and systems in a given scenario. This inductive human analysis facilitates better making sense of the illustrative descriptions generated by the LLM, as well as their critical evaluation and the identification of the most salient way of describing the themes that emerge from each scenario.